OpenStudy (anonymous):

Simplify:: 10+|10-2|+2^3 over 5^2-12. Help Please :)

jhonyy9 (jhonyy9):

can you makeing the calcule here step by step ?

jhonyy9 (jhonyy9):

I10-2I=

jhonyy9 (jhonyy9):

2^3 =

jhonyy9 (jhonyy9):

5^2 =

jhonyy9 (jhonyy9):

so you dont like cooperating nothing ?

jhonyy9 (jhonyy9):

than bye

OpenStudy (anonymous):

umm i was away from my desk... i know the answer is 2, just dont know how to get there. 10+8+8 over 25-12

hartnn (hartnn):

yes, now what is 10+8+8=?

hartnn (hartnn):

?

hartnn (hartnn):

still away from desk ?

OpenStudy (anonymous):

i solved it thanks... && yes i went to go pee.

hartnn (hartnn):

lol! welcome ^_^